# A Survey of Frontiers in LLM Reasoning: Inference Scaling, Learning to Reason, and Agentic Systems

### v3

@@ 6.2.1 Evaluating Reasoning
Evaluating reasoning beyond outcomes remains an open and challenging problem. Early approaches relied on human annotators to assess the quality of generated explanations (Camburu et al., 2018; Rajani et al., 2019), focusing on whether the reasoning could lead to the same predictions. To scale this idea, follow-up works (Wiegreffe et al., 2020; Hase et al., 2020) used trained models as simulators to evaluate the alignment between generated reasoning and final predictions. When human-annotated reasoning chains are available, some studies leverage traditional NLG metrics to measure overlap between human- and model-generated explanations (Clinciu et al., 2021). Others propose reasoning-specific metrics to assess aspects like coherency, redundancy, factuality (Golovneva et al., 2022), informativeness (Chen et al., 2022), robustness (Wang & Zhao, 2024), and contextual faithfulness (Ming et al., 2025). Under the LLM-as-Judge paradigm, recent works prompt powerful LLMs like GPT-4 to directly evaluate reasoning chains generated by other models (Hao et al., 2024a; Sun et al., 2024b). However, as reasoning tasks grow in complexity, evaluation becomes increasingly difficult, even for frontier models—if a model cannot perform a task, how can it judge if the task is done correctly? Thus, developing robust and accurate methods to evaluate reasoning beyond outcomes remains a significant and unresolved challenge.
@@ 6.2.2 Understanding Reasoning
Recent LLMs exhibit strong performance across diverse tasks, suggesting some level of reasoning capability. However, whether these skills are general and transferable or merely specialized for tasks encountered during pretraining remains an open and debated question. To address this, several empirical studies have sought to understand and enhance LLM capabilities across various reasoning forms: abstractive reasoning (Wu et al., 2024a; He & Lu, 2024), compositional reasoning (Bhargava & Ng, 2022; Li et al., 2024g), inductive reasoning (Yang et al., 2024f; Han et al., 2024b), abductive reasoning (Jung et al., 2022; Pareschi, 2023), deductive reasoning (Poesia et al., 2024; Seals & Shalin, 2024; Feng et al., 2024a), logical reasoning (Wan et al., 2024b; Han et al., 2024a; Xu et al., 2025c), commonsense reasoning (Lin et al., 2021; Liang et al., 2023a; Sun et al., 2024a), math reasoning (Ahn et al., 2024; Mirzadeh et al., 2025), and social reasoning (Gandhi et al., 2023). Notably, Arkoudas (2023) qualitatively evaluate GPT-4 on 21 diverse reasoning problems, concluding that despite occasional analytical success, GPT-4 remains incapable of true reasoning. Similarly, Wu et al. (2024a) empirically investigate abstractive reasoning and find that while LLMs achieve nontrivial performance on counterfactual tasks, their performance consistently degrades compared to default conditions, indicating reliance on narrow, non-transferable procedures. Mondorf & Plank (2024) provide a comprehensive survey on recent evaluations of LLM reasoning abilities.
@@
There is increasing interest in formal analyses, which use structured and logical proofs to systematically evaluate and improve the reasoning capabilities of LLMs. Han et al. (2022) introduce FOLIO, a dataset designed to assess models’ ability to derive correct conclusions from premises using first-order logic reasoning. Similarly, Saparov & He (2023) develop a benchmark evaluating LLMs on symbolic ontologies, revealing that models often struggle with proof planning and rely on knowledge retrieval rather than genuine reasoning. These findings highlight the potential of neurosymbolic methods to better understand LLM reasoning. Recent work also explores formal analysis techniques to enhance LLM reasoning. For instance, Pan et al. (2023) use LLMs to translate natural language problems into symbolic formulations, which are then processed by deterministic symbolic solvers for inference. (Li et al., 2025b) demonstrate the promise of leveraging LLMs’ symbolic reasoning for mathematical problem-solving. Other studies focus on domain-specific reasoning: Fang et al. (2024a) propose an LLM-based agent for text-based games, designed to tackle symbolic challenges and achieve in-game objectives, while Nahid & Rafiei (2024) introduce a framework to enhance LLMs’ symbolic reasoning by normalizing web tables. These studies reveal LLMs’ limitations in structured reasoning while emphasizing the value of integrating formal analysis to strengthen their capabilities.
@@
The body of work exploring the theoretical insights into CoT mechanisms remains relatively limited, with most studies focusing on the expressiveness of LLMs when using CoT. A pioneering study by Feng et al. (2023a) investigates LLMs with CoT for solving mathematical and decision-making problems. Using circuit complexity theory (Arora & Barak, 2009), they demonstrate that bounded-depth transformers cannot solve basic arithmetic or equation tasks unless the model size grows super-polynomially. In contrast, they prove that constant-size models can solve these tasks, along with a wide range of decision-making problems such as Dynamic Programming, by generating CoT derivations in a common mathematical language. Li et al. (2024h) extend these findings, providing a tighter upper bound on the expressiveness of constant-depth transformers with CoT. However, these studies do not explore how the length of a CoT affects model reasoning power. To address this gap, Merrill & Sabharwal (2024) find that a logarithmic number of intermediate steps (relative to input length) offers only marginal gains over standard transformers, while a linear number of steps under the assumption of projected pre-norm (a slight generalization of standard pre-norm) enables the recognition of all regular languages. Furthermore, polynomially many steps, combined with generalized pre-norm, allow transformers to recognize exactly the class of polynomial-time solvable problems.

### v4

@@ 6.2.1 Evaluating Reasoning
Evaluating reasoning beyond outcomes remains an open and challenging problem. Early approaches relied on human annotators to assess the quality of generated explanations (Camburu et al., 2018; Rajani et al., 2019), focusing on whether the reasoning could lead to the same predictions. To scale this idea, follow-up works (Wiegreffe et al., 2020; Hase et al., 2020) used trained models as simulators to evaluate the alignment between generated reasoning and final predictions. When human-annotated reasoning chains are available, some studies leverage traditional NLG metrics to measure overlap between human- and model-generated explanations (Clinciu et al., 2021). Others propose reasoning-specific metrics to assess aspects like coherency, redundancy, factuality (Golovneva et al., 2022), informativeness (Chen et al., 2022), robustness (Wang and Zhao, 2024), and contextual faithfulness (Ming et al., 2025). Under the LLM-as-Judge paradigm, recent works prompt powerful LLMs like GPT-4 to directly evaluate reasoning chains generated by other models (Hao et al., 2024a; Sun et al., 2024b). However, as reasoning tasks grow in complexity, evaluation becomes increasingly difficult, even for frontier models—if a model cannot perform a task, how can it judge if the task is done correctly? Thus, developing robust and accurate methods to evaluate reasoning beyond outcomes remains a significant and unresolved challenge.
@@ 6.2.2 Understanding Reasoning
Recent LLMs exhibit strong performance across diverse tasks, suggesting some level of reasoning capability. However, whether these skills are general and transferable or merely specialized for tasks encountered during pretraining remains an open and debated question. To address this, several empirical studies have sought to understand and enhance LLM capabilities across various reasoning forms: abstractive reasoning (Wu et al., 2024a; He and Lu, 2024), compositional reasoning (Bhargava and Ng, 2022; Li et al., 2024g), inductive reasoning (Yang et al., 2024e; Han et al., 2024b), abductive reasoning (Jung et al., 2022; Pareschi, 2023), deductive reasoning (Poesia et al., 2024; Seals and Shalin, 2024; Feng et al., 2024a), logical reasoning (Wan et al., 2024b; Han et al., 2024a; Xu et al., 2025c), commonsense reasoning (Lin et al., 2021; Liang et al., 2023a; Sun et al., 2024a), math reasoning (Ahn et al., 2024; Mirzadeh et al., 2025), and social reasoning (Gandhi et al., 2023). Notably, Arkoudas (2023) qualitatively evaluate GPT-4 on 21 diverse reasoning problems, concluding that despite occasional analytical success, GPT-4 remains incapable of true reasoning. Similarly, Wu et al. (2024a) empirically investigate abstractive reasoning and find that while LLMs achieve nontrivial performance on counterfactual tasks, their performance consistently degrades compared to default conditions, indicating reliance on narrow, non-transferable procedures. Mondorf and Plank (2024) provide a comprehensive survey on recent evaluations of LLM reasoning abilities.
@@
There is increasing interest in formal analyses, which use structured and logical proofs to systematically evaluate and improve the reasoning capabilities of LLMs. Han et al. (2022) introduce FOLIO, a dataset designed to assess models’ ability to derive correct conclusions from premises using first-order logic reasoning. Similarly, Saparov and He (2023) develop a benchmark evaluating LLMs on symbolic ontologies, revealing that models often struggle with proof planning and rely on knowledge retrieval rather than genuine reasoning. These findings highlight the potential of neurosymbolic methods to better understand LLM reasoning. Recent work also explores formal analysis techniques to enhance LLM reasoning. For instance, Pan et al. (2023) use LLMs to translate natural language problems into symbolic formulations, which are then processed by deterministic symbolic solvers for inference. (Li et al., 2025b) demonstrate the promise of leveraging LLMs’ symbolic reasoning for mathematical problem-solving. Other studies focus on domain-specific reasoning: Fang et al. (2024a) propose an LLM-based agent for text-based games, designed to tackle symbolic challenges and achieve in-game objectives, while Nahid and Rafiei (2024) introduce a framework to enhance LLMs’ symbolic reasoning by normalizing web tables. These studies reveal LLMs’ limitations in structured reasoning while emphasizing the value of integrating formal analysis to strengthen their capabilities.
@@
The body of work exploring the theoretical insights into CoT mechanisms remains relatively limited, with most studies focusing on the expressiveness of LLMs when using CoT. A pioneering study by Feng et al. (2023a) investigates LLMs with CoT for solving mathematical and decision-making problems. Using circuit complexity theory (Arora and Barak, 2009), they demonstrate that bounded-depth transformers cannot solve basic arithmetic or equation tasks unless the model size grows super-polynomially. In contrast, they prove that constant-size models can solve these tasks, along with a wide range of decision-making problems such as Dynamic Programming, by generating CoT derivations in a common mathematical language. Li et al. (2024h) extend these findings, providing a tighter upper bound on the expressiveness of constant-depth transformers with CoT. However, these studies do not explore how the length of a CoT affects model reasoning power. To address this gap, Merrill and Sabharwal (2024) find that a logarithmic number of intermediate steps (relative to input length) offers only marginal gains over standard transformers, while a linear number of steps under the assumption of projected pre-norm (a slight generalization of standard pre-norm) enables the recognition of all regular languages. Furthermore, polynomially many steps, combined with generalized pre-norm, allow transformers to recognize exactly the class of polynomial-time solvable problems.